OpenStudy (anonymous):

CHECK MY HOMEWORK PROBLEM PLEASE. 3C - 6D = 12 7C + 5D = 9 7(3C - 6D = 12)= 7(12) 3 (7C + 5D = 9)= 3(9) 21C - 42D = 84 21C - 15D = 27 21C - 42D = 84 -21C - 15D = -27 21C - 42D = 84 -21C - 15D = -27 _________________ 57D = 57 D = 57/-57= -1 3C - 6D = 12 3C - 6(-1)= 12 3C + 6 = 12 3C = 12 - 6 3C = 6 C = 6/3 = 2

OpenStudy (anonymous):

check your answers: 3(2)-6(-1)=12? 6+6=12? true 7c+5d=9? 7(2)+5(-1)=9? 14-5=9? true your answers are good

OpenStudy (ash2326):

@JohannaG You have done a great work. Thanks @Dockworker for explaining how to check the answers for correctness.:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Thank you guys:-)